# Neyman-Pearson multiclass classification under label noise via empirical likelihood

### 23 Mar 2026 v1

@@ Abstract
In many classification problems, the costs of misclassifying observations from different classes can be highly unequal. The Neyman–Pearson multiclass classification (NPMC) framework addresses this issue by minimizing a weighted misclassification risk while imposing upper bounds on class-specific error probabilities. Existing NPMC methods typically assume that training labels are correctly observed. In practice, however, labels are often corrupted due to measurement error or annotation, and the effect of such label noise on NPMC procedures remains largely unexplored. We study the NPMC problem when only noisy labels are available in the training data. We propose an empirical likelihood (EL)–based method that relates the distributions of noisy and true labels through an exponential tilting density ratio model. The resulting maximum EL estimators recover the class proportions and posterior probabilities of the clean labels required for error control. We establish consistency, asymptotic normality, and optimal convergence rates for these estimators. Under mild conditions, the resulting classifier satisfies NP oracle inequalities with respect to the true labels asymptotically. An expectation–maximization algorithm computes the maximum EL estimators. Simulations show that the proposed method performs comparably to the oracle classifier under clean labels and substantially improves over procedures that ignore label noise.
@@ 1 Introduction
In many real-world machine learning applications, especially in high-stakes domains such as medical diagnosis and financial fraud detection, the consequences of different classification errors are highly asymmetric. For instance, misclassifying a malignant tumor as benign can be far more consequential than the opposite error (Bokhari et al., 2021), and approving a fraudulent loan carries greater risk than rejecting a legitimate one. Traditional classifiers, which typically minimize the overall misclassification rate and treat all errors equally, do not account for such asymmetries.
@@
The Neyman–Pearson multiclass classification (NPMC) framework addresses this limitation by explicitly constraining error rates for selected classes, thereby prioritizing safety and reliability over simple accuracy. Specifically, let $(X,Y)$ be a pair of random variables where $X\in{\mathcal{X}}\subset{\mathbb{R}}^{p}$ denotes the feature vector and $Y\in[K]=\{0,1,\dots,K-1\}$ is the class label. For each $k$, let $P_{k}^{*}(A):={\mathbb{P}}(X\in A|Y=k)$ for any Borel set $A\subset{\mathcal{X}}$, and denote $X^{k}\triangleq X|(Y=k)\sim P_{k}^{*}$. The NPMC problem [^1] The NPMC problem reduces to classical overall misclassification error minimization when $\rho_{k}=w_{k}^{*}={\mathbb{P}}(Y=k)$ for all $k$ and $\mathcal{S}$$\varnothing$. seeks a classifier $\phi:{\mathcal{X}}\to[K]$ that solves:
@@
$\begin{split}\min_{\phi}&\quad{\mathcal{L}}(\phi)=\sum_{k=0}^{K-1}\rho_{k}P_{k}^{*}(\{\phi(X)\neq k\})\\ \text{subject to}&\quad P_{k}^{*}(\{\phi(X)\neq k\})\leq\alpha_{k},~\forall~k\in{\mathcal{S}}\subset[K],\end{split}$ (1)
where $\rho_{k}$, $\alpha_{k}$, and $\mathcal{S}$} are pre-specified weights, target error levels, and class indices with constrained errors, respectively.
When $K=2$, problem (1) reduces to the classical Neyman–Pearson binary classification task: minimizing the type II error while controlling the type I error. This binary setting has been extensively studied; see, for example, Scott and Nowak (2005); Scott (2007); Rigollet and Tong (2011); Tong (2013); Zhao et al. (2016); Tong et al. (2018, 2020); Li et al. (2020); Xia et al. (2021); Wang et al. (2024). In contrast, the general multiclass case ($K\geq 3$) remains relatively underexplored. Extending binary NP classification to the multiclass setting is non-trivial, in part because problem (1) may be infeasible for certain specifications of $\rho_{k}$ and $\alpha_{k}$. Recently, Tian and Feng (2025) proposed a dual-based algorithm that checks asymptotic feasibility and provides an efficient solver when strong duality holds. Their estimator further satisfies an NP oracle inequality, which compares the learned classifier $\widehat{\phi}$ to the oracle classifier $\phi^{*}$ that satisfies the error constraints:
$P_{k}^{*}(\{\widehat{\phi}(X)\neq k\})\leq P_{k}^{*}(\{\phi^{*}(X)\neq k\})+O_{p}(\varepsilon(n)),~\forall~k\in{\mathcal{S}}$ (2)
where $\varepsilon(n)$ vanishes as the sample size $n$ increases.
Despite its appeal, the NPMC framework fundamentally relies on clean, correctly labeled training data. This assumption is often violated in practice due to measurement error, annotator inconsistency, or adversarial noise. Empirical studies suggest that roughly 5% of labels in real-world datasets are incorrect (Yao et al., 2023). Label noise is present even in standard benchmarks: for instance, Northcutt et al. (2021) find mislabeling in datasets ranging from MNIST to ImageNet. In domain-specific datasets such as CheXpert (Irvin et al., 2019), ambiguities in radiology reports can also result in erroneous labels. More recently, in applications involving large language models, stochastic variability in model outputs has been shown to induce label noise relative to an underlying latent ground truth, further illustrating the pervasiveness of noisy labels in modern machine learning systems (Zhang and Martinez, 2025).
It is therefore natural to ask: What are the consequences of label noise for the NPMC problem? In the binary case, Yao et al. (2023) showed that ignoring label noise can lead to overly conservative type I error and thus an inflated type II error. Their approach, however, requires prior knowledge of the noise transition (confusion) matrix—the conditional probability of a noisy label given the true label—which is rarely available in practice. Moreover, extending such results to the multiclass setting is also non-trivial. More broadly, it remains unclear how label noise impacts the general NPMC framework and how it can be properly accounted for, a gap that this paper aims to fill.
Concretely, suppose we observe $n$ independently and identically distributed (i.i.d.) samples $\{(X_{i},Y_{i},\widetilde{Y}_{i})\}_{i=1}^{n}\sim P^{*}_{X,Y,\widetilde{Y}}$, where $Y_{i}$ denotes the unknown true label and $\widetilde{Y}_{i}$ is a noisy version of $Y_{i}$. In practice, we only observe the noisy training set ${\mathcal{D}}^{\text{train}}=\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}$, along with $m$ additional i.i.d. feature-only samples ${\mathcal{D}}^{\mathrm{test}}={X_{i}}{i=n+1}^{n+m}$ from the same distribution.
Our goal is to use only the noisy ${\mathcal{D}}^{\text{train}}$ to learn a classifier $\widehat{\phi}$ such that its predictions on ${\mathcal{D}}^{\text{test}}$ satisfy the constraints in (1) with respect to the (unobserved) true labels $\{Y_{i}\}_{i=n+1}^{n+m}$.
We propose a novel empirical likelihood (EL) framework for solving the NPMC problem under label noise, without requiring prior knowledge of the noise transition matrix. Our method captures the relationship between the true and noisy label distributions through an exponential tilting density ratio model (DRM). By estimating the DRM parameters, we recover the key quantities–true class proportions and posterior probabilities of the clean labels–needed for error control under clean labels, which allows us to directly apply the clean-label error control procedure of Tian and Feng (2025) to the noisy-label setting. Computationally, we treat the true labels as latent variables and develop an efficient expectation–maximization algorithm to maximize the profile empirical likelihood and consistently estimate these quantities.
The proposed method has several key advantages. First, it eliminates the need for any prior knowledge or bounds on the noise parameters, enhancing practical feasibility. Second, the resulting maximum empirical likelihood estimators are statistically consistent and asymptotically normal with optimal convergence rates. Third, under mild assumptions, we provide rigorous theoretical guarantees showing that the final classifier satisfies NP oracle inequalities. Fourth, simulation studies demonstrate that our method achieves near-oracle performance—closely matching the error control of classifiers trained on clean labels—while significantly outperforming naive approaches that ignore label noise. Finally, our framework naturally accommodates both binary and multiclass classification, thereby filling a notable gap in the existing literature.
@@
The rest of the paper is organized as follows. Section 2 introduces the problem setup and the relationship between noisy and clean labels. Section 3 presents the proposed EL-based estimator, its asymptotic normality, and the EM algorithm. Section 4 studies the binary NP problem, including the classifier, oracle guarantees, and empirical results. Section 5 extends the framework to the general NPMC setting. Section 6 presents real-data results, and Section 7 concludes.
@@ 2 Problem setting
To enable valid statistical inference, we must formalize the relationship between latent $Y$ and observed $\widetilde{Y}$. Following established work (Angluin and Laird, 1988; Yao et al., 2023; Sesia et al., 2024; Bortolotti et al., 2025), we adopt the standard instance independence assumption:
@@
This assumption captures the common scenario as shown in Sesia et al. (2024, Appendix A.3) where label noise depends primarily on class membership rather than specific feature values, as frequently observed in real-world settings like crowdsourcing (Snow et al., 2008). Beyond its practical relevance, this assumption provides crucial tractability by establishing a clear relationship between the contaminated distribution $\widetilde{P}_{k}^{*}$ of $\widetilde{X}^{k}$ and the true conditional distribution $P_{k}^{*}$ of $X^{k}$, as formalized in the proposition below.
@@
$\text{(Marginal)}\quad w_{k}^{*}=\sum_{l\in[K]}M_{lk}^{*}\widetilde{w}_{l}^{*},\quad\widetilde{w}_{l}^{*}=\sum_{k\in[K]}T_{lk}^{*}w_{k}^{*},$ (3)
$\text{(Conditional)}\quad\widetilde{P}_{l}^{*}=\sum_{k\in[K]}M_{lk}^{*}P_{k}^{*},\quad P_{k}^{*}=\sum_{l\in[K]}T_{lk}^{*}\widetilde{P}_{l}^{*},$ (4)
$\text{(Posterior)}\quad\widetilde{\pi}_{l}^{*}(x)=\sum_{k\in[K]}T_{lk}^{*}\pi_{k}^{*}(x).\quad\quad\quad\quad\quad$ (5)
[…]
Remark 2.1.
[…]
Based on these relationships, a naive two-step approach to estimating $\pi_{k}^{*}(x)$ via (5) would require prior knowledge of the transition matrix $\bm{T}$$^{*}$, which is typically unavailable in practice. Instead, we propose a joint estimation procedure that simultaneously learns $\bm{T}$$^{*}$ and $\pi_{k}^{*}(x)$ by directly leveraging these distributional relationships, which we detail in the next section.

### 21 Apr 2026 v2

@@ Abstract
In many classification problems, misclassification costs are highly asymmetric, while training labels are often corrupted due to measurement error, annotator variability, or adversarial noise. The Neyman–Pearson multiclass classification (NPMC) framework addresses such asymmetry by controlling class-specific errors, but existing methods assume that training labels are correctly observed. To our knowledge, no existing approach handles NPMC under label noise in the multiclass setting, and the only binary method requires prior knowledge of the noise mechanism. A fundamental difficulty is that, without structural assumptions, noisy-label models are non-identifiable: distinct combinations of class-conditional distributions and noise mechanisms can induce the same observed distribution, preventing recovery of the quantities required for error control. We show that the exponential tilting density ratio model restores identifiability, and leverage this structure to develop an empirical likelihood approach for NPMC with noisy labels. The proposed method jointly estimates clean-label class proportions, posterior probabilities, and the noise mechanism from noisy data, without requiring prior knowledge of the confusion matrix. An expectation–maximization algorithm enables efficient computation. The resulting estimators are $\sqrt{n}$-consistent and asymptotically normal, and the induced classifiers satisfy Neyman–Pearson oracle inequalities in both binary and multiclass settings. Simulation and real-data experiments demonstrate near-oracle performance.
@@ 1 Introduction
1.1 Label noise is pervasive in real-world applications
In supervised learning, label noise refers to discrepancies between observed labels and the (unobserved) ground truth. Such contamination is ubiquitous in practice. Empirical studies suggest that roughly 5% of labels in real-world datasets are incorrect, and even widely used benchmarks such as MNIST and ImageNet contain non-negligible annotation errors (Northcutt et al., , 2021). In domain-specific settings, for example in CheXpert (Irvin et al., , 2019), ambiguities in radiology reports give rise to label errors that are difficult to quantify yet unavoidable. More recently, outputs from LLMs have also been shown to induce label noise (Zhang and Martinez, , 2025).
Label noise may arise from sources including measurement error, annotator variability, and adversarial corruption. Its impact on statistical learning procedures is well documented (Frénay and Verleysen, , 2013): contaminated labels can distort decision boundaries, bias parameter estimates, and lead to overly complex models whose apparent performance, when evaluated against noisy labels, substantially overstates their accuracy on clean data (Northcutt et al., , 2021). A substantial body of work (Natarajan et al., , 2013; Patrini et al., , 2017; Han et al., 2018a, ; Han et al., 2018b, ; Wu et al., , 2021; Wang et al., , 2022) addresses label noise in general classification settings. However, they largely focus on improving overall classification accuracy, and does not account for settings where different types of errors must be controlled in an asymmetric manner, as we discuss below.
1.2 Neyman–Pearson multiclass classification under clean labeled data
@@
In many high-stakes applications, minimizing overall misclassification error is not the appropriate objective. The consequences of different types of errors are often highly asymmetric: for example, misclassifying a malignant tumor as benign is far more consequential than the reverse error (Bokhari et al., , 2021), and approving a fraudulent loan carries greater risk than rejecting a legitimate one. In such settings, the goal is not to minimize average error, but to control errors for specific high-risk classes while reducing errors elsewhere.
@@
The Neyman–Pearson (NP) classification framework formalizes this principle by seeking a classifier that minimizes a weighted misclassification risk subject to explicit upper bounds on class-specific error probabilities. Let $(X,Y)$ be a pair of random variables, where $X\in{\mathcal{X}}\subset{\mathbb{R}}^{p}$ denotes the feature vector and $Y\in[K]=\{0,1,\dots,K-1\}$ the class label. The Neyman–Pearson multiclass classification (NPMC) problem [^1] The NPMC problem reduces to classical overall misclassification error minimization when $\rho_{k}=w_{k}^{*}={\mathbb{P}}(Y=k)$ for all $k$ and $\mathcal{S}$$\varnothing$. seeks a classifier $\phi:{\mathcal{X}}\to[K]$ that solves:
@@
$\begin{split}\min_{\phi}&\qquad{\mathcal{L}}(\phi)=\sum_{k=0}^{K-1}\rho_{k}P_{k}^{*}(\{\phi(X)\neq k\})\\ \text{subject to}&\qquad P_{k}^{*}(\{\phi(X)\neq k\})\leq\alpha_{k},~\forall~k\in{\mathcal{S}}\subset[K],\end{split}$ (1)
where $\rho_{k}$, $\alpha_{k}$, and $\mathcal{S}$} are pre-specified weights, target error levels, and constrained class indices respectively, and $P_{k}^{*}(A):={\mathbb{P}}(X\in A|Y=k)$ for any Borel set $A\subset{\mathcal{X}}$. When $K=2$, problem (1) reduces to the classical NP binary classification task of minimizing the type II error subject to a constraint on the type I error, a setting that has been extensively studied (Scott and Nowak, , 2005; Scott, , 2007; Rigollet and Tong, , 2011; Tong, , 2013; Zhao et al., , 2016; Tong et al., , 2018, 2020; Li et al., , 2020; Xia et al., , 2021; Wang et al., , 2024).
The general multiclass case ($K\geq 3$) is considerably more challenging and has received comparatively less attention. In particular, problem (1) may be infeasible for certain specifications of $\rho_{k},\alpha_{k}$}. Tian and Feng, (2025) addressed this issue via a Lagrangian dual formulation, showing that the dual objective is concave irrespective of the convexity of the primal problem, thereby enabling efficient optimization using standard methods. Their framework further provides a feasibility check and establishes NP oracle inequalities for the resulting classifier. A key feature of this approach is its plug-in structure: given consistent estimators of the class proportions $w_{k}^{*}={\mathbb{P}}(Y=k)$} and the class-posterior probabilities $\pi_{k}^{*}(x)={\mathbb{P}}(Y=k|X=x)$} under clean labels, the optimal classifier can be obtained by substitution into the dual solution. Despite these appealing properties, the framework explicitly assumes that training labels are correctly observed—an assumption that is routinely violated in precisely the applications as we show in Section 1.1.
1.3 Challenges of NPMC under noisy labeled data
The plug-in structure of Tian and Feng, (2025) suggests, at least in principle, a natural route to handling label noise: rather than redesigning the solver, one may instead supply consistent estimators of $w_{k}^{*}$ and $\pi_{k}^{*}(x)$. However, constructing such estimators from noisy labeled data alone is far from straightforward and constitutes the central statistical challenge addressed in this paper.
When only noisy labels $\widetilde{Y}$ are observed, naive estimators target the contaminated quantities $\{\widetilde{w}_{k}^{*}={\mathbb{P}}(\widetilde{Y}=k)\}$ and $\{\widetilde{\pi}_{k}^{*}(x)={\mathbb{P}}(\widetilde{Y}=k|X=x)\}$ rather than their clean-label counterparts. Substituting these into the NPMC solver yields classifiers calibrated to the noisy label distribution, which in turn leads to systematic violations of the error constraints with respect to the true labels. For instance, Yao et al., (2023) shows that ignoring label noise results in overly conservative type I error control and, consequently, inflated type II error under the binary NP setting. A natural strategy is to apply existing noisy-label methods as a preprocessing step before solving the NPMC problem. However, this approach is inadequate for our purposes. Loss-correction methods such as Natarajan et al., (2013) and Patrini et al., (2017) are designed to recover the Bayes-optimal classifier for overall accuracy under noisy data, rather than to provide consistent estimators of the class proportions $w_{k}^{*}$ and posterior probabilities $\pi_{k}^{*}(x)$ required by the NPMC solver. Moreover, they rely on strong assumptions, including knowledge of the confusion matrix ${\bm{M}}^{*}=\{{\mathbb{P}}(Y=k|\widetilde{Y}=l)\}$ and, in the case of Patrini et al., (2017), the existence of anchor points. The confident learning approach of Northcutt et al., (2021) likewise targets improvements in overall classification performance and does not yield the estimators needed for NPMC.
To the best of our knowledge, Yao et al., (2023) is the only work that addresses asymmetric error control under label noise. Their analysis is restricted to the binary NP setting, where they propose a noise-adjusted NP umbrella algorithm. However, their method requires prior knowledge of noise confusion matrix $\bm{M}$$^{*}$. Such information is rarely available in practice, and the method does not provide a way to estimate these quantities. In the absence of such information, one may instead impose bounds on these quantities; however, loose bounds can lead to performance comparable to that of naive methods that ignore label noise. Moreover, their approach does not extend to the multiclass setting, leaving the multiclass problem under label noise unaddressed.
A common feature of these existing approaches is their reliance on knowledge of the noise confusion matrix $\bm{M}$$^{*}$, which is typically unavailable in practice. A natural alternative is therefore to estimate this quantity from the observed data. However, this leads to a more fundamental difficulty: without additional structural assumptions, the noisy-label model is non-identifiable. In particular, distinct combinations of class-conditional distributions $\{P_{k}^{*}\}$ and confusion matrices can induce the same observed data distribution, so that the quantities required for NPMC error control cannot, in general, be recovered. This non-identifiability is not merely a theoretical concern but arises in concrete ways that preclude naive estimation. For example, convex perturbations of the class-conditional distributions can be absorbed into a modified noise mechanism without altering the observed distribution (Example 2.1). These observations indicate that the problem is inherently ill-posed without suitable structural constraints, motivating the approach in this paper.
1.4 Our contributions and overview
We address the Neyman–Pearson multiclass classification problem under label noise by resolving a fundamental identifiability barrier. We begin by formalizing the non-identifiability of noisy-label classification through concrete examples, showing that the problem is ill-posed without additional structural assumptions (Example 2.1). To overcome this difficulty, we introduce an exponential tilting density ratio model (DRM), under which the class-conditional distributions are modeled relative to a nonparametric reference distribution. We show that, in the absence of feature collinearity, this structure is sufficient to uniquely identify the confusion matrix, the clean-label class proportions $\{w_{k}^{*}\}$, and the posterior probabilities $\{\pi_{k}^{*}(x)\}$ from noisy observations alone, and we formalize this via a strict identifiability result for both the DRM parameters and the reference distribution (Theorem 3.1). To the best of our knowledge, this result has no direct analogue in the existing DRM literature and forms the foundation of our approach.
Building on this identification result, we develop an empirical likelihood (EL) framework for estimation and inference in the NPMC problem under the standard instance-independent noise assumption $\widetilde{Y}\perp\!\!\!\!\perp X|Y$. The parameters $\{w_{k}^{*}\}$ and $\{\pi_{k}^{*}(x)\}$ are jointly estimated via the maximum empirical likelihood estimators (MELEs) from noisy training data, without requiring knowledge of the noise transition matrix. We show that these estimators are $\sqrt{n}$-consistent and asymptotically normal (Theorem 3.2).
For efficient computation, we treat the true labels as latent variables and develop an expectation-maximization (EM) algorithm to maximize the profile empirical likelihood. The resulting procedure is straightforward to implement, as the M-step reduces to a weighted multinomial logistic regression. We further establish monotonic ascent of the objective function (Proposition 3.1), which guarantees convergence of the algorithm.
Finally, we show that the resulting plug-in classifiers satisfy Neyman–Pearson oracle inequalities with respect to the true labels, in both binary (Theorem 4.1) and multiclass settings (Theorem 5.1). In particular, our method achieves valid error control without requiring prior knowledge of the noise transition matrix and provides, to the best of our knowledge, the first such guarantees for multiclass NPMC under label noise. Empirical results on simulated and real datasets demonstrate near-oracle performance and substantial improvements over existing approaches.
@@ 1.4 Our contributions and overview
The rest of the paper is organized as follows. Section 2 introduces the problem setup and the relationship between noisy and clean labels. Section 3 presents the EL-based estimator, its asymptotic properties, and the EM algorithm. Section 4 studies the binary NP problem, including the classifier, oracle guarantees, and simulations. Section 5 extends the framework to the multiclass setting. Section 6 presents real-data experiments, and Section 7 concludes.
@@ 2 Problem setting
Suppose we observe $n$ independent and identically distributed samples $\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}$, where $X_{i}$ denotes the feature vector, $Y_{i}$ the unobserved true label, and $\widetilde{Y}_{i}$ its noisy observed version. To enable valid statistical inference, we must formalize the relationship between latent $Y$ and observed $\widetilde{Y}$. Following established work (Angluin and Laird, , 1988; Yao et al., , 2023; Sesia et al., , 2024; Bortolotti et al., , 2025), we adopt the standard instance independence assumption:
@@
This assumption captures the common scenario as shown in Sesia et al., (2024, Appendix A.3) where label noise depends primarily on class membership rather than specific feature values, as frequently observed in real-world settings like crowdsourcing (Snow et al., , 2008). Beyond its practical relevance, this assumption provides crucial tractability by establishing a clear relationship between the contaminated distribution $\widetilde{P}_{k}^{*}$ of $\widetilde{X}^{k}\triangleq X|(\widetilde{Y}=k)$ and the true conditional distribution $P_{k}^{*}$ of $X^{k}\triangleq X|(Y=k)$, as formalized in the proposition below.
@@
$\text{(Marginal)}\qquad w_{k}^{*}=\sum_{l\in[K]}M_{lk}^{*}\widetilde{w}_{l}^{*},\qquad\widetilde{w}_{l}^{*}=\sum_{k\in[K]}T_{lk}^{*}w_{k}^{*},$ (2)
$\text{(Conditional)}\qquad\widetilde{P}_{l}^{*}=\sum_{k\in[K]}M_{lk}^{*}P_{k}^{*},\qquad P_{k}^{*}=\sum_{l\in[K]}T_{lk}^{*}\widetilde{P}_{l}^{*},$ (3)
$\text{(Posterior)}\qquad\widetilde{\pi}_{l}^{*}(x)=\sum_{k\in[K]}T_{lk}^{*}\pi_{k}^{*}(x).\qquad\qquad\qquad\qquad\qquad$ (4)
The proof is deferred to Appendix C.1. Based on these relationships, a naive two-step approach to estimating $\pi_{k}^{*}(x)$ via (4) would require prior knowledge of the transition matrix $\bm{T}$$^{*}$, which is typically unavailable in practice. In the absence of such knowledge, we emphasize that, without additional structural assumptions, the model is fundamentally unidentifiable, beyond the usual label permutation ambiguity encountered in mixture models. In particular, multiple distinct parameterizations can induce exactly the same observed distribution. We illustrate this phenomenon through the following example.
Example 2.1 (Algebraic unidentifiability of true distributions).
If $P_{0}^{*}(x)$ and $P_{1}^{*}(x)$ are unconstrained, the model is unidentifiable. For any small constant $\epsilon\in(0,1)$, we can define a new, "contaminated" candidate distribution for class 1: $P_{1}^{**}(x)\triangleq(1-\epsilon)P_{1}^{*}(x)+\epsilon P_{0}^{*}(x).$ Because $P_{1}^{**}(x)$ is a convex combination of two valid probability density functions, it is also a valid density function. We can rewrite the true $P_{1}^{*}(x)$ in terms of this new distribution:
$\widetilde{P}_{0}(x)=(1-\rho_{0}^{*})P_{0}^{*}(x)+\rho_{1}^{*}\left[\frac{1}{1-\epsilon}P_{1}^{**}(x)-\frac{\epsilon}{1-\epsilon}P_{0}^{*}(x)\right]=\left(1-\rho_{0}^{*}-\frac{\epsilon\rho_{1}^{*}}{1-\epsilon}\right)P_{0}^{*}(x)+\left(\frac{\rho_{1}^{*}}{1-\epsilon}\right)P_{1}^{**}(x).$}
This example demonstrates that simultaneous estimation of $\bm{T}$$^{*}$ and $\{\pi_{k}^{*}(x)\}$ requires additional structural assumptions on these distributions, which are introduced in the next section.